### v1

# Decomposition theorems for unmatchable pairs in groups and field extensions

Abstract.
A theory of matchings for finite subsets of an abelian group, introduced in connection with a conjecture of Wakeford on canonical forms for symmetric tensors, has since been extended to the setting of field extensions and to that of matroids. Earlier approaches have produced numerous criteria for matchability and unmatchability, but have offered little structural insight. In this paper, we develop parallel structure theorems which characterize unmatchable pairs in both abelian groups and field extensions. Our framework reveals analogous obstructions to matchability: nearly periodic decompositions of sets in the group setting correspond to decompositions of subspaces involving translates of simple extensions in the linear setting. This perspective not only recovers previously known results through short proofs, but also leads to new matching criteria and guarantees the existence of nontrivial unmatchable pairs.
1Department of Mathematics, Clayton State University, 2000 Clayton State Boulevard, Lake City, Georgia 92093, USA. maliabadi@clayton.edu.
2Department of Mathematics, Long Island University, 720 Northern Blvd, Brookville, New York 11548, USA. Jozsef.Losonczy@liu.edu.
Corresponding Author.
Keywords and phrases. Chowla sets and subspaces, Field extension, Matching property, Nearly periodic decomposition
2020 Mathematics Subject Classification. Primary: 05D15; Secondary: 11B75; 12F99.
1. Introduction
Overview. Let $G$ be an abelian group with operation written multiplicatively, and let $A,B\subseteq G$ be nonempty finite sets. A matching from $A$ to $B$ is a bijection $f:A\to B$ such that $af(a)\notin A$ for all $a\in A$. This notion was introduced by Fan and the second author of the present paper [9] in a work on Wakeford’s conjecture concerning canonical forms for symmetric tensors [14]. The approach taken in [9] involved the study of a specific class of matchings, called acyclic matchings, for certain pairs of subsets of $\mathbb{Z}^{n}$. When such matchings exist for all pairs $(A,B)$ satisfying obvious necessary conditions, the ambient group is said to have the acyclic matching property. This property was established for $\mathbb{Z}^{n}$ in [6], extended to torsion-free abelian groups in [12], and fully characterized for all abelian groups in [4].
Further progress has included broader results in abelian groups [1], an extension to arbitrary groups [7], enumerative results using the isoperimetric method [10], a linear formulation over field extensions [8], and a matroidal analogue [5]. For an application of matchings in the context of combinatorial number theory, see [11].
Recently, in [3], the present authors established characterization theorems for matchable subsets in abelian groups and matchable subspaces over field extensions using Dyson transforms. These results removed a previous reliance on a diverse collection of number-theoretic inequalities and their linear analogues. The present paper continues that program by giving new structural characterizations of unmatchable sets and subspaces and exploring their implications, which include existence criteria for nontrivial unmatchable pairs.
Organization of the paper. In Sections 1.1 and 1.2, we recall background information on matchings in abelian groups and field extensions. Section 2 establishes a structure theorem for unmatchable pairs $(A,B)$ of finite subsets of any abelian group (Theorem 2.3), exhibiting a nearly periodic decomposition for such pairs, and then develops consequences, e.g., the symmetric case $A=B$, a Chowla-type criterion, and inheritance of unmatchability to suitable quotients. Section 3 presents a linear analogue for field extensions: any unmatchable pair of subspaces decomposes as a direct sum built from translates of simple extensions (Theorem 3.3). Several corollaries follow, including results on self-matching and Chowla subspaces. Section 4 describes some directions for further work.
1.1. Matchings in abelian groups.
We begin with some terminology and notation. For any positive integer $n$, we write $[n]$ for $\{1,\ldots,n\}$. Let $A$ and $B$ be nonempty finite sets of the same cardinality and let $\Delta$ be a subset of $A\times B$. A bijective mapping $f:A\longrightarrow B$ is called a matching of $\Delta$ if $(a,f(a))\in\Delta$ for all $a\in A$. We have a certain algebraic context in mind for the sets $A$, $B$, and $\Delta$. Let $G$ be an abelian group and let $A$ and $B$ be nonempty finite subsets of $G$ such that $|A|=|B|$. Define
$\Delta=\{(a,b)\in A\times B:ab\notin A\}.$
We will always assume that $\Delta$ is as above. A matching of $\Delta$ is then a bijection $f:A\longrightarrow B$ satisfying $af(a)\notin A$ for all $a\in A$. In the case where there exists at least one matching of $\Delta$, we say that the pair $(A,B)$ is matchable; otherwise, we call $(A,B)$ unmatchable.
In order for $(A,B)$ to be matchable, it is clearly necessary that $1\notin B$. We say that the group $G$ has the matching property if this condition is also sufficient. The following results were established in the abelian group setting in [12]:
  • For a finite nonempty subset $A$ of $G$, the pair $(A,A)$ is matchable if and only if $1\notin A$.
  • $G$ has the matching property if and only if $G$ is torsion-free or cyclic of prime order.
Our main objective in the first part of this paper will be to provide a structural decomposition of unmatchable sets in $G$ (Theorem 2.3). Several applications of this result will then be presented.
1.2. Matching subspaces in a field extension.
For any subset $S$ of a vector space $V$, we let $\langle S\rangle$ denote the subspace of $V$ spanned by $S$. If $S=\{x_{1},\ldots,x_{n}\}$, we also write $\langle x_{1},\ldots,x_{n}\rangle$ for this subspace. Given a field extension $K\subseteq L$ and $K$-subspaces AA and $B$ of $L$, we use $AB$ to denote the Minkowski product of $A$ and $B$:
$AB=\{ab:a\in A,\ b\in B\}.$
In [8], Eliahou and Lecouvey introduced a notion of matchability for subspaces in a field extension. We give their definition below.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be $n$-dimensional $K$-subspaces of $L$ with $n>0$. An ordered basis $\mathcal{A}=\{a_{1},\ldots,a_{n}\}$ of $A$ is said to be matched to an ordered basis $\mathcal{B}=\{b_{1},\ldots,b_{n}\}$ of $B$ if, for each $i\in[n]$,
$a_{i}^{-1}A\cap B\subseteq\langle b_{1},\ldots,b_{i-1},b_{i+1},\ldots,b_{n}\rangle.$
Note that when the above condition holds, we have $a_{i}b_{i}\notin\mathcal{A}$ for all $i$, so the correspondence $a_{i}\mapsto b_{i}$ defines a matching, in the group-theoretic sense, from $\mathcal{A}$ to $\mathcal{B}$ in the multiplicative group of $L$.
We say that the subspace $A$ is matched to $B$ if every ordered basis of $A$ can be matched to some ordered basis of $B$. In this case, we also say that the pair $(A,B)$ is matchable; otherwise, we call $(A,B)$ unmatchable.
A necessary condition for $A$ to be matched to $B$ is $1\notin B$. This is discussed in detail in [8].
A field extension $K\subsetneq L$ is said to have the linear matching property if, for every pair of finite-dimensional $K$-subspaces AA and $B$ of $L$ with $\dim A=\dim B>0$ and $1\notin B$, the subspace $A$ is matched to $B$.
Eliahou and Lecouvey established the following fundamental results [8]:
  • A subspace $A$ is matched to itself if and only if $1\notin A$.
  • A field extension $K\subsetneq L$ has the linear matching property if and only if $L$ contains no nontrivial proper finite-dimensional extension over $K$.
In Section 3, we will present a decomposition theorem for unmatchable pairs of subspaces (Theorem 3.3). Several consequences of this result will then be developed. We will also give a brief summary of the correspondence between relevant concepts in the group and linear settings (see Table 1).
2. Structure theorem for unmatchable pairs in abelian groups
The theorem below is one of the main results of [3]. It can be used to recover, efficiently and without recourse to supplementary ideas, all previously known results about matchings in abelian groups. In this section, it will be used to derive a structure theorem for unmatchable pairs $(A,B)$.
Theorem 2.1.
Let $A$ and $B$ be nonempty finite subsets of an abelian group $G$, with $|A|=|B|$ and $1\notin B$. Then $(A,B)$ is matchable if and only if, for every pair of nonempty subsets $S\subseteq A$ and $R\subseteq B\cup\{1\}$ with $SR=S$, we have
$|S|\leq|B\setminus R|.$
For any subset $X$ of an abelian group $G$, we will write $\langle X\rangle$ for the subgroup of $G$ generated by $X$.
The condition $SR=S$ in Theorem 2.1 can be formulated in a different way, as described in the following proposition. For a proof, see [3].
Proposition 2.2.
Let $G$ be an abelian group and let $S$ and $R$ be nonempty finite subsets of $G$. Then $SR=S$ if and only if $S$ is a union of cosets of $\langle R\rangle$.
Below, we state and prove our structure theorem for unmatchable pairs $(A,B)$ in an abelian group. It will be used to derive new criteria for the existence and nonexistence of matchings, and it will serve as the basis for a new and concise proof of a known result concerning Chowla sets. An analogous result in the setting of field extensions will be presented in Section 3.
Theorem 2.3.
Let $A$ and $B$ be nonempty finite subsets of an abelian group $G$, with $|A|=|B|$. Then the pair $(A,B)$ is unmatchable if and only if there exists a nonempty subset $R$ of $B$ such that $A$ and $B$ can be expressed as disjoint unions as follows:
$A$ $=S\cup Y,$
$B$ $=R\cup Z,$
where $S$ is a union of cosets of $\langle R\rangle$, and YYsatisfies $|Y|<|R|$.
Note: The conditions on the sets in the above decomposition ensure that $S\neq\emptyset$.
Proof.
Assume that the pair $(A,B)$ is unmatchable. Observe that if $1\in B$, then we can simply take $S=A$, $Y=\emptyset$, $R=\{1\}$, and $Z=B\setminus R$.
Suppose $1\notin B$. Then, by Theorem 2.1, there exist sets $S$ and $R$ satisfying the following conditions:
  • (a)
    $\emptyset\neq S\subseteq A$ and $\emptyset\neq R\subseteq B\cup\{1\}$,
  • (b)
    $SR=S$,
  • (c)
    $|S|>|B\setminus R|$.
We see from (b) and Proposition 2.2 that $S$ is a union of cosets of $\langle R\rangle$. Note that $R\neq\{1\}$, by (c). Moreover, we may assume that $1\notin R$, since otherwise we can replace $R$ with $R\setminus\{1\}$, and (a)–(c) will still hold.
Thus $R\subseteq B$, and we can now rewrite (c) as
|R|>|B||S|.|R|>|B|-|S|.
Take $Y=A\setminus S$ and $Z=B\setminus R$. Since $|A|=|B|$ and $S\subseteq A$, the above inequality shows that $|Y|<|R|$, giving us the desired decomposition.
Conversely, assume that $A$ and $B$ can be written as in the statement. If $B$ contains $1$, then clearly $(A,B)$ is unmatchable, so assume $1\notin B$. The sets $S$ and $R$ satisfy $\emptyset\neq S\subseteq A$, $\emptyset\neq R\subseteq B$, and $SR=S$. Furthermore, the inequality $|Y|<|R|$ implies $|S|+|R|>|S|+|Y|=|A|=|B|$, hence $|S|>|B\setminus R|$. Therefore, by Theorem 2.1, the pair $(A,B)$ is unmatchable. ∎
Let $G$ be an abelian group and let $H$ be a subgroup. A subset $X$ of $G$ is called $H$-periodic if $X$ is a nonempty union of cosets of $H$. Any decomposition as in Theorem 2.3,
$A=S\cup Y,\qquad B=R\cup Z,$
will be called a nearly periodic decomposition for the unmatchable pair $(A,B)$, with $\langle R\rangle$-periodic set $S$ and remainder $Y.$
Corollary 2.4.
Let $A$ and $B$ be nonempty finite subsets of an abelian group $G$, with $|A|=|B|$ and $1\notin A$. Assume that for each nonempty subset $R$ of $B$, we have $|\langle R\rangle\cap A|\geq|R|$. Then the pair $(A,B)$ is matchable.
Proof.
Assume that $(A,B)$ is unmatchable. Then the pair $(A,B)$ has a nearly periodic decomposition
$A=S\cup Y,\qquad B=R\cup Z,$
with $R,S,Y,Z$ as described in Theorem 2.3. Note that $\langle R\rangle\cap S=\emptyset$, since $1\notin A$ and $S$ is a union of cosets of $\langle R\rangle$. Therefore
$|\langle R\rangle\cap A|=|\langle R\rangle\cap Y|\leq|Y|<|R|.$
We can use the above corollary to recover Theorem 2.1 from [12].
Corollary 2.5.
Let $A$ be a nonempty finite subset of an abelian group, with $1\notin A$. Then $(A,A)$ is matchable.
Proof.
Take $A=B$ in Corollary 2.4. ∎
A nonempty finite subset $B$ of a group $G$ is called a Chowla set if $o(x)>|B|$ for every $x\in B$, where $o(x)$ denotes the order of $x$. In [10], Hamidoune showed that if $B$ is a Chowla set, then the pair $(A,B)$ is matchable for any set $A$ with $|A|=|B|$. Using the structure theorem above, it is possible to give a very short proof of this fact when $G$ is abelian.
Corollary 2.6.
Let $A$ and $B$ be nonempty finite subsets of an abelian group $G$, with $|A|=|B|$. Assume that $B$ is a Chowla set. Then the pair $(A,B)$ is matchable.
Proof.
By the Chowla assumption, for each nonempty subset $R$ of $B$, we have $|\langle R\rangle|>|B|=|A|$, and so it is impossible for $A$ to be decomposed as in the statement of Theorem 2.3. ∎
If $G$ is a group having at least one finite nontrivial proper subgroup, we denote by $n_{0}(G)$ the smallest size of such a subgroup. The next result is new.
Corollary 2.7.
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$ and $n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$. Then there exist nonempty finite subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair.
Proof.
We begin by writing $n=q\cdot n_{0}(G)+r$, where $q$ and $r$ are integers such that $q\geq 1$ and $0\leq r<n_{0}(G)$. Let $x$ be an element of $G$ of order $n_{0}(G)$. Define $R$ by
$R=\langle x\rangle\setminus\{1\}=\{x,x^{2},\ldots,x^{n_{0}(G)-1}\},$
and let $S$ be the union of any $q$ distinct cosets of $\langle R\rangle=\langle x\rangle$. Let $Y$ be any subset of $G\setminus S$ of size $r$, and let $Z$ be any subset of $G\setminus\langle x\rangle$ of size $n-n_{0}(G)+1$. Finally, define $A$ and $B$ by
$A=S\cup Y,\qquad B=R\cup Z.$
Then $1\notin B$ and $|A|=|B|=n$. Also, since $n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$, it follows that
$|Y|=r<n_{0}(G)-1=|R|.$
Thus, the above is a nearly periodic decomposition for $(A,B)$. By Theorem 2.3, this pair is unmatchable. ∎
Remark 2.8.
It turns out that when
$n\equiv n_{0}(G)-1\pmod{n_{0}(G)}$,
nontrivial unmatchable pairs $(A,B)$ with $|A|=|B|=n$ may or may not exist. For instance, in the additive group $\mathbb{Z}/4\mathbb{Z}$, every pair $(A,B)$ of $3$-element subsets with $\bar{0}\notin B$ is matchable. On the other hand, in $\mathbb{Z}/12\mathbb{Z}$, the $5$-element subsets
$A=\{\bar{0},\bar{1},\bar{3},\bar{6},\bar{9}\},\qquad B=\{\bar{1},\bar{2},\bar{3},\bar{6},\bar{9}\}$
(with $\bar{0}\notin B$) form an unmatchable pair.
Although the condition
$n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$
in Corollary 2.7 cannot be omitted, the result below establishes a necessary and sufficient condition for the existence of an unmatchable pair $(A,B)$ with $1\notin B$ in all cases, including when $n$ lies in the congruence class of $n_{0}(G)-1\pmod{n_{0}(G)}$.
Theorem 2.9.
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$. Then there exist nonempty finite subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair if and only if there is a finite subgroup $H$ of $G$ with $|H|\leq n$ and $|H|\nmid(n+1)$.
Proof.
Suppose an unmatchable pair $(A,B)$ as described in the statement exists. Then, by Theorem 2.3, there is a nonempty subset $R$ of $B$ such that $A$ and $B$ can be expressed as disjoint unions
$A=S\cup Y,\qquad B=R\cup Z,$
where $S$ is a union of cosets of $\langle R\rangle$, and the remainder $Y$ satisfies $|Y|<|R|$. We will show that $\langle R\rangle$ can serve as the desired subgroup $H$ of $G$.
First note that $\langle R\rangle$ is finite, since $S$ is. Let $m=|\langle R\rangle|$ and write $|S|=mq$, where $q\geq 1$. Then
$n=|S|+|Y|=mq+|Y|.$
Clearly $m\leq mq\leq n$. We also have $|Y|<|R|\leq m-1$, as $1\notin R$, and so
$mq\leq n\leq mq+m-2.$
Hence
$mq<mq+1\leq n+1\leq mq+m-1<m(q+1).$
From this we see that $m\nmid(n+1)$, as desired.
Conversely, assume that $G$ has a subgroup $H$ with $|H|\leq n$ and $|H|\nmid(n+1)$. Let $m=|H|$ and note that $m\geq n_{0}(G)>1$. We can write
$n=mq+r$
for some integers $q\geq 1$ and $0\leq r\leq m-2$. Set $R=H\setminus\{1\}$ and let $S$ be a union of $q$ distinct cosets of $\langle R\rangle$. Then $|S|=mq$. Let $Y\subseteq G\setminus S$ with $|Y|=r$, and let $Z\subseteq G\setminus(R\cup\{1\})$ with |Z|=nm+1|Z|=n-m+1. Finally, define
$A=S\cup Y,\qquad B=R\cup Z.$
Then $1\notin B$ and $|A|=|B|=n$. Further, all of the conditions in Theorem 2.3 are satisfied, implying that $(A,B)$ is unmatchable. ∎
Remark 2.10.
When $n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$, the condition of Theorem 2.9 is automatically met for any $H$ with $|H|=n_{0}(G)$. In this way, we retrieve Corollary 2.7.
In the following corollary, we establish a connection between unmatchable sets in a given abelian group $G$ and their images in a quotient of $G$.
Corollary 2.11.
Let $G$ be an abelian group and let $(A,B)$ be a pair of nonempty finite subsets with $|A|=|B|$. Suppose $(A,B)$ is unmatchable, as witnessed by the decomposition given in Theorem 2.3:
$A=S\cup Y,$
where $S$ is a union of cosets of the subgroup $\langle R\rangle$ for some nonempty $R\subseteq B$, and $|Y|<|R|$. Suppose $H$ is a subgroup of $G$ satisfying $H\cap(SS^{-1}\cup RR^{-1})=\{1\}$, and let $\pi:G\to G/H$ be the corresponding quotient map. Then the projected pair $(\pi(A),\pi(B))$ is unmatchable in $G/H$.
Proof.
First notice that if $|\pi(A)|\neq|\pi(B)|$, then the pair $(\pi(A),\pi(B))$ is unmatchable, and we are done. So suppose $|\pi(A)|=|\pi(B)|$. Define $Y^{\prime}=\pi(A)\setminus\pi(S)$ and $Z^{\prime}=\pi(B)\setminus\pi(R)$. Then we obtain the decompositions
$\pi(A)=\pi(S)\cup Y^{\prime}\quad\text{and}\quad\pi(B)=\pi(R)\cup Z^{\prime}.$
These satisfy the conditions of Theorem 2.3, since
  • $\pi(S)$ is a nonempty union of cosets of $\langle\pi(R)\rangle$,
  • $|Y^{\prime}|<|\pi(R)|$.
To see why the latter holds, observe first that
$|Y^{\prime}|=|\pi(A)|-|\pi(S)|\leq|A|-|\pi(S)|.$
Now, $|\pi(S)|=|S|$ and $|\pi(R)|=|R|$ since $H\cap(SS^{-1}\cup RR^{-1})=\{1\}$. Combining this with the above, we obtain
$|Y^{\prime}|\leq|A|-|S|=|Y|<|R|=|\pi(R)|.$
Therefore, $(\pi(A),\pi(B))$ is unmatchable in $G/H$. ∎
Remark 2.12.
The assumption $H\cap(SS^{-1}\cup RR^{-1})=\{1\}$ in Corollary 2.11 cannot be dropped. To see this, consider the additive group $G=\mathbb{Z}/8\mathbb{Z}$, the subgroup $H=\{\bar{0},\bar{4}\}=4\mathbb{Z}/8\mathbb{Z}$, and the corresponding quotient map
$\pi:G\longrightarrow G/H\;\cong\;\mathbb{Z}/4\mathbb{Z}.$
Define
$A=\{\bar{0},\bar{1},\bar{2},\bar{4},\bar{6}\},\qquad B=\{\bar{1},\bar{2},\bar{3},\bar{5},\bar{6}\}.$
The following decomposition of $A$ and $B$ satisfies the assumptions of Theorem 2.3:
$A=S\cup Y,\qquad B=R\cup Z,$
where $R=\{\bar{2},\bar{6}\}$, $S=\bar{0}+\langle R\rangle=\{\bar{0},\bar{2},\bar{4},\bar{6}\}$, $Y=\{\bar{1}\}$, and $Z=\{\bar{1},\bar{3},\bar{5}\}$. Hence $(A,B)$ is unmatchable in $\mathbb{Z}/8\mathbb{Z}$. Note that $H\cap((S-S)\cup(R-R))\neq\{\bar{0}\}$.
We now compute $\pi(A)$ and $\pi(B)$:
$\pi(A)=\{\bar{0},\bar{1},\bar{2}\}\subseteq\mathbb{Z}/4\mathbb{Z},\qquad\pi(B)=\{\bar{1},\bar{2},\bar{3}\}\subseteq\mathbb{Z}/4\mathbb{Z}.$
These are matchable: for instance, the bijection $\bar{0}\mapsto\bar{3}$, $\bar{1}\mapsto\bar{2}$, $\bar{2}\mapsto\bar{1}$ is a matching since each sum lands at $\bar{3}$, which lies outside $\pi(A)$.
Thus $(A,B)$ is unmatchable in $G$, but $(\pi(A),\pi(B))$ is matchable in $G/H$.
3. Structure theorem for unmatchable subspaces in a field extension
The following theorem, a main result of [3], provides a framework for studying matchable subspaces in field extensions. Using it, one can recover all known matching criteria in this setting. We will apply it to obtain a structure theorem for unmatchable pairs $(A,B)$.
Theorem 3.1.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$ and $1\notin B$. Then $A$ is matched to $B$ if and only if, for every pair of nonzero $K$-subspaces $S\subseteq A$ and $R\subseteq B\oplus K$ with $\langle SR\rangle=S$, we have
$\dim S\leq\dim\bigl(B/(R\cap B)\bigr).$
The proposition below explains, in large part, the significance of the condition $\langle SR\rangle=S$ in Theorem 3.1. A proof is given in [3].
Proposition 3.2.
Let $K\subseteq L$ be a field extension, let $n$ be a positive integer, let $S$ and $R$ be nonzero $K$-subspaces of $L$, and let $x\in R$. Assume that $\langle SR\rangle=S$ and $\dim S\leq n$. Then $x$ is algebraic over $K$ and in fact $K(x):K]\leq n$. Also, for each $a\in S$, we have $aK(x)\subseteq S$.
We next state and prove the principal result of this section, giving a structural decomposition for unmatchable pairs $(A,B)$ over a field extension $K\subsetneq L$. This theorem is a linear analog of Theorem 2.3. Here, translates of simple field extensions, i.e., $K$-subspaces of the form $aK(x)$ with $a,x\in L$, play a role similar to that of cosets in the group setting. We note that such subspaces were considered briefly in the matching context in [2].
Theorem 3.3.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$. Then the pair $(A,B)$ is unmatchable if and only if there exists a nonzero subspace $R\subseteq B$ such that $A$ and $B$ can be written as
$A$ $=S\oplus Y,$
$B$ $=R\oplus Z,$
where
$S=\sum_{\begin{subarray}{c}a\in S\end{subarray}}aK(x)$
for every $x\in R$, and $\dim Y<\dim R$.
Note: The conditions on the subspaces in the above decomposition ensure that $S$ is nonzero.
Proof.
Assume that the pair $(A,B)$ is unmatchable. If $1\in B$, define $S=A$, $Y=\{0\}$, $R=K$, and choose $Z$ to be a $K$-linear complement of $K$ in $B$ (so that $B=K\oplus Z$). The required conditions are then satisfied.
Now suppose $1\notin B$. By Theorem 3.1, there exist nonzero subspaces $S$ and $R$ satisfying the following:
  • (a)
    $S\subseteq A$ and $R\subseteq B\oplus K$,
  • (b)
    $\langle SR\rangle=S$,
  • (c)
    $\dim S>\dim\bigl(B/(R\cap B)\bigr)$.
By condition (b) and Proposition 3.2, we have $aK(x)\subseteq S$ for every $a\in S$ and $x\in R$, i.e., $S$ is a vector space over $K(x)$ for each $x\in R$. In other words,
$S=\sum_{a\in S}aK(x)$
for every $x\in R$. Note that $R\neq K$ by (c), and we may assume that $R\subseteq B$; otherwise, we can replace $R$ by $\pi_{B}(R)$, where $\pi_{B}$ is the projection map $B\oplus K\longrightarrow B$ along $K$, and conditions (a)–(c) will still hold. Thus, condition (c) gives us
$\dim R>\dim B-\dim S.$
Now let $Y$ be a complementary summand of $A$ relative to $S$, and $Z$ be a complementary summand of $B$ relative to $R$. Since $\dim A=\dim B=n$, it follows from the inequality above that $\dim Y<\dim R$, as desired.
Conversely, assume that $A$ and $B$ can be written as in the statement. If $1\in B$, then the pair $(A,B)$ is unmatchable, so suppose $1\notin B$. Then $S$ and $R$ are nonzero subspaces with $S\subseteq A$, $R\subseteq B$, and $\langle SR\rangle=S$. Furthermore, the inequality $\dim Y<\dim R$ implies
$\dim R+\dim S>\dim Y+\dim S=\dim A=\dim B.$
This yields
$\dim S>\dim B-\dim R=\dim B/R=\dim\bigl(B/(B\cap R)\bigr).$
Therefore, by Theorem 3.1, the pair $(A,B)$ is unmatchable. ∎
In the following table, we summarize the main information regarding unmatchable pairs in abelian groups and their linear analogues.
Group setting Linear setting
Objects $A,B\subseteq G$, $|A|=|B|<\infty$ KK–subspaces $A,B\subseteq L$, $\dim A=\dim B<\infty$
Decomposition $A=S\cup Y$, $B=R\cup Z$ (disjoint unions) $A=S\oplus Y$, $B=R\oplus Z$ (direct sums)
Structure of $S$ $S$ is a union of $\langle R\rangle$–cosets $S=\sum_{a\in S}aK(x)$ for all $x\in R$
Gap condition $|Y|<|R|$ $\dim Y<\dim R$
Approach Find $\emptyset\neq R\subseteq B$ with $S$ as above and $|Y|<|R|$ Find $\{0\}\neq R\subseteq B$ with $S$ as above and $\dim Y<\dim R$
Consequence $(A,B)$ is an unmatchable pair $(A,B)$ is an unmatchable pair
Table 1. Side-by-side dictionary for the decomposition of unmatchable pairs.
For our first application of Theorem 3.3, we give a new, short proof of Theorem 4.15 from [2].
Corollary 3.4.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$. Suppose that $A$ is matched to $B$. Then $\langle AB\rangle\neq A$.
Proof.
Assume that $\langle AB\rangle=A$. Set $S=A$, $R=B$, and $Y=Z=\{0\}$. By Proposition 3.2, we have
$S=\sum_{a\in S}aK(x)$
for all $x\in R$. It is easy to see that the sets $S,R,Y,Z$ satisfy all of the hypotheses of Theorem 3.3. Therefore AA is not matched to $B$. ∎
Let $K\subsetneq L$ be a field extension, and let $B$ be a $K$-subspace of $L$. We say that $B$ is a Chowla subspace if, for every $x\in B\setminus\{0\}$,
$K(x):K]\geq\dim B+1.$
In particular, if $B$ is a Chowla subspace, then $1\notin B$. The following result, originally proved in [3], is retrieved easily using Theorem 3.3.
Corollary 3.5.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two finite-dimensional $K$-subspaces of $L$, with $\dim A=\dim B>0$. Assume that $B$ is a Chowla subspace. Then $A$ is matched to $B$.
Proof.
For all $0\neq a\in A$ and $0\neq x\in B$, we have
$\dim aK(x)=[K(x):K]>\dim B=\dim A,$
by the Chowla assumption. Thus a decomposition of $A$ as in Theorem 3.3 is not possible. We conclude that $A$ is matched to $B$. ∎
Example 3.6.
Assume $K=\mathbb{Q}$ and $L=\mathbb{Q}(\sqrt[p]{2})$, where $p$ is a prime. Let $A$ be any $\mathbb{Q}$-subspace of $L$ of dimension p1p-1, and let $B$ be the trace-zero subspace of $\mathbb{Q}(\sqrt[p]{2})$ over $\mathbb{Q}$, which has dimension p1p-1. Then $B$ is a Chowla subspace. By Corollary 3.5, it follows that $A$ is matched to $B$.
The lemma below will help us to establish our next corollary. It will also be used in the proof of Theorem 3.12 later in this section.
Lemma 3.7.
Let $K\subseteq L$ be a field extension, and let $R$ and $S$ be $K$-subspaces of $L$ of finite dimension. Assume that $S$ is nonzero and $aK(x)\subseteq S$ for all $a\in S$ and $x\in R$. Then $aK(R)\subseteq S$ for all $a\in S$. Also, $K(R):K]<\infty$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$.
Proof.
Let $\{x_{1},\ldots,x_{t}\}$ be a $K$-basis for $R$, and let $i\in[t]$. By assumption, we can view $S$ as a vector space over $K(x_{i})$. Consider the equation
$\dim_{K}S=\dim_{K(x_{i})}S\cdot[K(x_{i}):K].$
Since $0<\dim_{K}S<\infty$, the above implies $K(x_{i}):K]<\infty$, and so $x_{i}$ is algebraic over $K$. We thus have $K(R)=K(x_{1},\ldots,x_{t})=K[x_{1},\ldots,x_{t}]$. It is now clear that $aK(R)\subseteq S$ for all $a\in S$, i.e., $S$ is a vector space over $K(R)$. In turn, this gives us
$\dim_{K}S=\dim_{K(R)}S\cdot[K(R):K],$
so that $K(R):K]<\infty$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$. ∎
The following result is new. It can be viewed as a linear analogue of Corollary 2.4.
Corollary 3.8.
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$ and $1\notin A$. Assume that for every nonzero $K$-subspace $R$ of $B$, we have
$\dim(K(R)\cap A)\geq\dim R.$
Then the pair $(A,B)$ is matchable.
Proof.
Suppose that the pair $(A,B)$ is unmatchable. Then, by Theorem 3.3, there exists a decomposition
$A=S\oplus Y,\qquad B=R\oplus Z,$
where $R$ is a nonzero subspace, $S=\sum_{a\in S}aK(x)$ for every $x\in R$, and $\dim Y<\dim R$. Hence, by Lemma 3.7,
$S=\sum_{a\in S}aK(R).$
Since we also have $1\notin A$, it follows that $K(R)\cap S=\{0\}$. Consider now the projection map $\pi:A\to Y$ along $S$, i.e., the $K$-linear map satisfying
$\ker\pi=S\qquad\text{and}\qquad\pi|_{Y}=\mathrm{id}_{Y}.$
Restrict $\pi$ to the subspace $T=K(R)\cap A$. Then
$\ker(\pi|_{T})=T\cap\ker\pi=(K(R)\cap A)\cap S=K(R)\cap S=\{0\}.$
Thus $\pi|_{T}$ is injective, and we obtain
$\dim(K(R)\cap A)=\dim T=\dim\pi(T)\leq\dim Y<\dim R.$
Using the above corollary, we can quickly obtain the commutative case of Theorem 2.8 in [8].
Corollary 3.9.
Let $K\subsetneq L$ be a field extension, and let $A$ be a nonzero finite-dimensional $K$-subspace of $L$. Then $A$ is matched to itself if and only if $1\notin A$.
Proof.
If $1\notin A$, we can take $A=B$ in Corollary 3.8. The converse follows immediately from the definition of matchable pair. ∎
For a field extension $K\subsetneq L$ with at least one nontrivial intermediate field of finite $K$-dimension, we define $n_{0}(K,L)$ to be the smallest degree of such an extension, i.e.,
$n_{0}(K,L)=\min\{[F:K]\mid K\subsetneq F\subseteq L,\;[F:K]<\infty\}.$
The following result, which can be viewed as a linear analogue of Corollary 2.7, guarantees the existence of nontrivial unmatchable pairs of sufficiently large dimension $n$, assuming that $n$ avoids a certain congruence class.
Corollary 3.10.
Let $K\subsetneq L$ be a field extension with at least one proper, nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that $n_{0}(K,L)\leq n<[L:K]$ and $n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$. Then there exist $K$-subspaces AA and $B$ of $L$ such that $1\notin B$, $\dim A=\dim B=n$, and the pair $(A,B)$ is unmatchable.
Proof.
Write $n=q\cdot n_{0}(K,L)+r$, where $q\geq 1$ and $0\leq r<n_{0}(K,L)$. Let $x\in L$ be such that
$K(x):K]=n_{0}(K,L).$
Define the $K$-subspace
$R=\langle x,x^{2},\dots,x^{n_{0}(K,L)-1}\rangle.$
Let $a_{1},\dots,a_{q}\in L$ be linearly independent over $K(x)$. Such elements exist since
$L:K(x)]>n/n_{0}(K,L)\geq q.$
Define
$S=\sum_{i=1}^{q}a_{i}K(x).$
The above sum is direct, hence
$\dim S=\sum_{i=1}^{q}\dim a_{i}K(x)=q\cdot[K(x):K]=q\cdot n_{0}(K,L).$
Observe that $S$ can be expressed as
$S=\sum_{a\in S}aK(x).$
Choose a $K$-subspace $Y\subseteq L$ such that $Y\cap S=\{0\}$ and $\dim Y=r$, and a $K$-subspace $Z\subseteq L$ such that $Z\cap K(x)=\{0\}$ and $\dim Z=n-n_{0}(K,L)+1$.
Now define
$A=S+Y,\qquad B=R+Z.$
These sums are direct, and since $n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$, we have
$\dim Y=r<n_{0}(K,L)-1=\dim R.$
Also, $\dim A=\dim B=n$, and $1\notin B$ since $1\notin R$ and $Z\cap K(x)=\{0\}$. Finally, for each $y\in R$, the field $K(y)$ is contained in $K(x)$, and so $S=\sum_{a\in S}aK(y)$. Therefore, by Theorem 3.3, the pair $(A,B)$ is unmatchable. ∎
Remark 3.11.
The condition
$n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$
in Corollary 3.10 cannot be omitted. For example, consider the field extension $\mathbb{Q}\subseteq\mathbb{Q}(\sqrt[4]{2})$. We claim that every pair $(A,B)$ of $3$-dimensional $\mathbb{Q}$-subspaces of $\mathbb{Q}(\sqrt[4]{2})$ with $1\notin B$ is matchable.
Suppose for a contradiction that some such pair is unmatchable. Then, by Theorem 3.3, there exists a nonzero subspace $R\subseteq B$ such that $(A,B)$ admits the decompositions
$A=S\oplus Y,\qquad B=R\oplus Z,$
where for every $x\in R$,
$S=\sum_{a\in S}a\,\mathbb{Q}(x),$
and $\dim Y<\dim R$. Note that $S\neq\{0\}$. Also, we see that $\dim\mathbb{Q}(x)\leq\dim S\leq 3$ for all $x\in R$.
Now, the only nontrivial, proper intermediate subfield of $\mathbb{Q}\subseteq\mathbb{Q}(\sqrt[4]{2})$ is $\mathbb{Q}(\sqrt{2})$. Therefore, since $\{0\}\neq R\subseteq B$ and $1\notin B$, we have $\mathbb{Q}(x)=\mathbb{Q}(\sqrt{2})$ for some (in fact, all) nonzero $x\in R$. Hence
$S=\sum_{a\in S}a\,\mathbb{Q}(\sqrt{2}).$
The above shows that $S$ is a vector space over $\mathbb{Q}(\sqrt{2})$. As such, its $\mathbb{Q}$-dimension must be even. This forces $\dim S=2$, $\dim Y=1$, and then $\dim R\geq 2$.
On the other hand, since $\mathbb{Q}(x)=\mathbb{Q}(\sqrt{2})$ for every nonzero $x\in R$, it follows that $R\subseteq\mathbb{Q}(\sqrt{2})$, so $\dim R\leq 2$. Hence $\dim R=2$, and we now have $R=\mathbb{Q}(\sqrt{2})$, which contradicts the assumption that $1\notin B$.
Therefore, the pair $(A,B)$ is matchable.
In Theorem 3.12 below, we provide a necessary and sufficient condition for the existence of a nontrivial unmatchable pair $(A,B)$ in all situations where $n_{0}(K,L)$ is defined and less than $[L:K]$, including when the dimension $n$ of $A$ and $B$ satisfies
$n\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$.
This result is a linear analogue of Theorem 2.9.
Theorem 3.12.
Let $K\subsetneq L$ be a field extension with at least one proper, nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that
$n_{0}(K,L)\leq n<[L:K].$
Then there exist $n$-dimensional $K$-subspaces $A,B\subseteq L$ with $1\notin B$ such that $(A,B)$ is an unmatchable pair if and only if there exists an intermediate extension $K\subsetneq F\subseteq L$ with $F:K]\leq n$ and $F:K]\nmid(n+1)$.
Proof.
Assume there exist $n$-dimensional $K$-subspaces $A,B\subseteq L$ such that $1\notin B$ and $(A,B)$ is unmatchable. Applying Theorem 3.3, we obtain subspaces $R,S,Y,Z$ as described therein. Observe that since
$S=\sum_{a\in S}aK(x)$
for every $x\in R$, it follows that $S$ is a vector space over each field $K(x)$ for $x\in R$. Hence, by Lemma 3.7, $S$ is a vector space over $K(R)$ and $K(R):K]<\infty$.
Let $F=K(R)$ and $m=[F:K]$. Again by Lemma 3.7, the $K$-dimension of $S$ must be a positive multiple of $m$. We now show that $F$ can serve as the desired intermediate field of $K\subsetneq L$. Write $\dim S=mq$ where $q\geq 1$. Then
$n=\dim A=\dim S+\dim Y=mq+\dim Y.$
Clearly $m\leq mq\leq n$. Observe also that since $\dim Y<\dim R$, $R\subseteq K(R)=F$, and $R\cap K=\{0\}$, we have $\dim Y\leq m-2$. This gives us
$mq\leq n\leq mq+m-2.$
Hence
$mq<mq+1\leq n+1\leq mq+m-1<m(q+1),$
so that $m\nmid(n+1)$, as desired.
Conversely, assume that there exists an intermediate extension $K\subsetneq F\subseteq L$ with $F:K]\leq n$ and $F:K]\nmid(n+1)$. Let $m=[F:K]\geq 2$ and observe that we can write
$n=mq+r,$
where $q\geq 1$ and $0\leq r\leq m-2$.
Next we define $R$ and $S$. Choose a subspace $R\subseteq F$ with $\dim R=m-1$ and $R\cap K=\{0\}$. Note that $L:F]>n/m\geq q$, and let $a_{1},\dots,a_{q}\in L$ be linearly independent over $F$. Define
$S=\sum_{i=1}^{q}a_{i}F.$
This sum is direct. Let $Y$ be a $K$-subspace of $L$ with $Y\cap S=\{0\}$ and $\dim Y=r$. Also, let $Z$ be a $K$-subspace of $L$ with $Z\cap(R+K)=\{0\}$ and $\dim Z=n-m+1$.
Finally, define
$A=S+Y,\qquad B=R+Z.$
Both of these sums are direct, and $1\notin B$. Moreover, the conditions of Theorem 3.3 are all satisfied, so that the pair $(A,B)$ is unmatchable. ∎
Remark 3.13.
We can recover Corollary 3.10 from Theorem 3.12 by observing that, when $n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$, the condition of Theorem 3.12 is automatically satisfied for any intermediate field $F$ with $F:K]=n_{0}(K,L)$.
The corollary below addresses the special case of finite fields.
Corollary 3.14.
Consider the field extension $\mathbb{F}_{q}\subsetneq\mathbb{F}_{q^{m}}$, where $q$ is a prime power and $m$ is a composite number. Let $t$ be the smallest prime divisor of $m$, and let $n$ be an integer with $t\leq n<m$. Then there exist unmatchable subspaces $A,B\subseteq\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ with $\dim A=\dim B=n$ and $1\notin B$ if and only if there exists a positive divisor $d$ of $m$ such that $d\leq n$ and $d\nmid(n+1)$.
Proof.
This follows immediately from Theorem 3.12, since the intermediate fields of the extension $\mathbb{F}_{q}\subseteq\mathbb{F}_{q^{m}}$ are precisely the fields of the form $\mathbb{F}_{q^{d}}$, where $d$ is a positive divisor of $m$, and $\mathbb{F}_{q^{d}}:\mathbb{F}_{q}]=d$. ∎
4. Future directions
  1. (1)
    We believe that a non-abelian version of Theorem 2.3 is plausible. The natural question is: what replaces the nearly periodic decomposition in this setting, and can obstructions be described in terms of coset-like structures or subgroup actions?
  2. (2)
    A natural next step would be to investigate pairs $(A,B)$ which are partially matchable, that is, matchable up to a prescribed defect. Some useful guidance may be provided by the “defect versions” of Hall’s marriage theorem and Rado’s theorem (see [13]).
  3. (3)
    One of the core ideas underlying our structural theorems is an implementation of Dyson’s $e$-transform. To what extent can this tool (and its linear analogue) be adapted to produce finer structural decompositions, beyond the binary distinction between matchable and unmatchable?
  4. (4)
    From an algorithmic perspective, given subsets (or subspaces) $A$ and $B$, can one efficiently detect whether they admit a nearly periodic decomposition (or simple-extension decomposition in the linear case)? What is the computational complexity of deciding unmatchability in these contexts?